OpenStudy (trumpdonald):

help! if you dont my mom will kill me! What is the equation in point-slope form of the line that passes through the points (7, 5) and (−4, −1)

OpenStudy (trumpdonald):

y-1= 6/11 (x-4)

OpenStudy (mrs.ambrose614):

Did I help you trumpdonald

OpenStudy (wolf1728):

First find the slope Slope = m = (Y2 -Y1) ÷ (X2 -X1) Slope = m = (-1-5) / (-4 -7) Slope = m = -6 / -11 = 6/11 Now we must put numbers in this equation: y = mx + b We can calculate "b" from this equation: b = y -mx Let's take one of the points (7,5) and put it into the equation b = 5 -(6/11 * 7) b = 5 -(42/11) b = 13/11 So, y = 6/11 x + 13/11